# You Only Sample Once: Taming One-Step Text-to-Image Synthesis by Self-Cooperative Diffusion GANs

### v5

@@ Abstract
Recently, some works have tried to combine diffusion and Generative Adversarial Networks (GANs) to alleviate the computational cost of the iterative denoising inference in Diffusion Models (DMs). However, existing works in this line suffer from either training instability and mode collapse or subpar one-step generation learning efficiency. To address these issues, we introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis with high training stability and mode coverage. Specifically, we smooth the adversarial divergence by the denoising generator itself, performing self-cooperative learning. We show that our method can serve as a one-step generation model training from scratch with competitive performance. Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models by several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the flawed noise scheduler). Experimental results show that YOSO achieves the state-of-the-art one-step generation performance even with Low-Rank Adaptation (LoRA) fine-tuning. In particular, we show that the YOSO-PixArt-$\alpha$ can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without extra explicit training, requiring only ~10 A800 days for fine-tuning. Our code is provided at https://github.com/Luo-Yihong/YOSO.
@@ 3.1 Our Design
Hence given the generator $G_{\theta}$ fixed, if we increase the information in ${\mathbf{x}}_{t}$, it is supposed that we can get a better distribution. In other words, it is highly likely that $p^{(t_{k})}_{\theta}({\mathbf{x}})$ is superior to $p^{(t)}_{\theta}({\mathbf{x}})$, where $t_{k}=\max\{t-k,0\}<t$. Motivated by the cooperative approach (Xie et al., 2018; 2021; 2022; Hill et al., 2022) which uses an MCMC revised version of model distribution to learn the generator, we suggest to use $p^{(t_{k})}_{\theta}({\mathbf{x}})$ as ground truth to learn $p^{(t)}_{\theta}({\mathbf{x}})$, which constructs the following training objective:
$\min_{\theta}\mathcal{L}_{\theta}\triangleq\mathbb{E}_{t}\mathbb{E}_{q({\mathbf{x}})q({\mathbf{x}}_{t}|{\mathbf{x}})}\lambda_{t}\times||G_{\theta}({\mathbf{x}}_{t},t)-{\mathbf{x}}||_{2}^{2}+\mathbb{E}_{t}(D_{\mathrm{adv}}(p^{(t_{k})}_{\theta}(\texttt{sg}({\mathbf{x}}))||p^{(t)}_{\theta}({\mathbf{x}})),$ (3)
where $\texttt{sg}[\cdot]$ denotes stop-gradient operator and the second term is named as cooperative adversarial loss. This training objective can be regarded as a self-cooperative approach, since the ‘revised’ samples from $p^{(t_{k})}_{\theta}({\mathbf{x}})$ and samples from $p^{(t)}_{\theta}({\mathbf{x}})$ are generated by the same network $G_{\theta}$. Note that we only replace data distribution as $p^{(t_{k})}_{\theta}({\mathbf{x}})$ in the adversarial divergence for smoothing the learning objective, as recent work (Luo et al., 2023d) has found that it is beneficial to learn the generator with a mix of real data and revised data.
@@ 6.1 Text-to-Image Generation
Experiment Setting. We initialize the GAN generator by pre-trained PixArt-$\alpha$ which is a diffusion transformer with 0.6B parameters. For the GAN discriminator, we construct the latent discriminator by pre-trained SD 1.5 using the proposed construction introduced in Sec. 5.1. We switch the pre-trained PixArt-$\alpha$ to v-prediction by the proposed technique introduced in Sec. 5.1, followed by training on the JourneyDB dataset (Pan et al., 2023) with resizing as 512 resolution. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We observe the training convergence fast, requiring only 30k iterations and around 10 A800 days to be trained. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We also conduct experiments on fine-tuning SD 1.5 via LoRA to show the effectiveness of the proposed YOSO, with a batch-size of 128 and a constant learning rate of 2e-5.
Figure 4: Qualitative comparisons of YOSO against competing methods.
@@
Table 2: Comparison of machine metrics on text-to-to-image generation across state-of-the-art methods. HFL denotes human feedback learning which might hack the machine metrics.
@@ 7 Related Works
Few-Step Text-to-Image Generation. LCM (Luo et al., 2023a) adapts consistency distillation (Song et al., 2023) for stable diffusion (SD) and enables image generation in 4 steps with acceptable quality. InstaFlow (Liu et al., 2023) adapts rectified flows (Liu et al., 2022; Liu, 2022) for SD, facilitating the generation of text-to-image in a mere single step. Despite this, the fidelity of the generated images is still poor. Some works (Yin et al., 2023) adapts variation score distillation (Wang et al., 2023) and diff-instruct (Luo et al., 2023c) for SD, which requires training an additional DMs for the generator distribution, akin to the discriminator in GANs. We found models trained in this way seem to have serious mode collapse issues (see Fig. 5). Recently, combining DMs and GANs for one-step text-to-image generation has been explored. UFOGen (Xu et al., 2023c) extends DDGANs (Xiao et al., 2022) and SSIDMs (Xu et al., 2023b) to SD by modifying the computation of reconstruction loss from corrupted samples to clean samples. However, it still performs adversarial matching using corrupted samples. ADD (Sauer et al., 2023c) introduces a one-step text-to-image generation based on SD. It follows earlier research (Sauer et al., 2023a) by employing a pre-trained image encoder, DINOv2 (Oquab et al., 2024), as the backbone of the discriminator to accelerate the training. However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands. Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training. This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training. SDXL-Lightning (Lin et al., 2024b) combines progressive distillation and adversarial training in a progressive framework. Recently, a concurrent work Hyper-SD (Ren et al., 2024) combines consistency distillation and adversarial training in a progressive framework, while our work can be trained end-to-end. However, their work performs adversarial point-level matching via consistency loss, and they still perform adversarial training via injecting noise. This results in less effective one-step learning, thus they rely on human feedback learning for achieving better performance. In contrast, we perform the adversarial training at the distribution level by replacing real data with self-generated data to smooth the adversarial divergence, which not only can stabilize the training without injecting noise but also form effective one-step generation learning. Moreover, compared with above mentioned diffusion-GAN hybrid models, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them. Additionally, we extend our method not only to SD but also to PixArt-$\alpha$ (Chen et al., 2024) which is based on diffusion transformer (Peebles & Xie, 2022). This demonstrates the wide application of our proposed YOSO.

### v6

@@ Abstract
Recently, some works have tried to combine diffusion and Generative Adversarial Networks (GANs) to alleviate the computational cost of the iterative denoising inference in Diffusion Models (DMs). However, existing works in this line suffer from either training instability and mode collapse or subpar one-step generation learning efficiency. To address these issues, we introduce YOSO, a novel generative model designed for rapid, scalable, and high-fidelity one-step image synthesis with high training stability and mode coverage. Specifically, we smooth the adversarial divergence by the denoising generator itself, performing self-cooperative learning. We show that our method can serve as a one-step generation model training from scratch with competitive performance. Moreover, we extend our YOSO to one-step text-to-image generation based on pre-trained models by several effective training techniques (i.e., latent perceptual loss and latent discriminator for efficient training along with the latent DMs; the informative prior initialization (IPI), and the quick adaption stage for fixing the flawed noise scheduler). Experimental results show that YOSO achieves the state-of-the-art one-step generation performance even with Low-Rank Adaptation (LoRA) fine-tuning. In particular, we show that the YOSO-PixArt-$\alpha$ can generate images in one step trained on 512 resolution, with the capability of adapting to 1024 resolution without extra explicit training, requiring only ~10 A800 days for fine-tuning. Our code is available at: https://github.com/Luo-Yihong/YOSO.
@@ 3.1 Our Design
Hence given the generator $G_{\theta}$ fixed, if we increase the information in ${\mathbf{x}}_{t}$, it is supposed that we can get a better distribution. In other words, it is highly likely that $p^{(t_{k})}_{\theta}({\mathbf{x}})$ is superior to $p^{(t)}_{\theta}({\mathbf{x}})$, where $t_{k}=\max\{t-k,0\}<t$. Motivated by the cooperative approach (Xie et al., 2018; 2021; 2022; Luo et al., 2024a) which uses an MCMC revised version of model distribution to learn the generator, we suggest to use $p^{(t_{k})}_{\theta}({\mathbf{x}})$ as ground truth to learn $p^{(t)}_{\theta}({\mathbf{x}})$, which constructs the following training objective:
$\min_{\theta}\mathcal{L}_{\theta}\triangleq\mathbb{E}_{t}\mathbb{E}_{q({\mathbf{x}})q({\mathbf{x}}_{t}|{\mathbf{x}})}\lambda_{t}||G_{\theta}({\mathbf{x}}_{t},t)-{\mathbf{x}}||_{2}^{2}+\mathbb{E}_{t}(D_{\mathrm{adv}}(p^{(t_{k})}_{\theta}(\texttt{sg}({\mathbf{x}}))||p^{(t)}_{\theta}({\mathbf{x}})),$ (3)
where $\texttt{sg}[\cdot]$ denotes stop-gradient operator and the second term is named as cooperative adversarial loss. This training objective can be regarded as a self-cooperative approach, since the ‘revised’ samples from $p^{(t_{k})}_{\theta}({\mathbf{x}})$ and samples from $p^{(t)}_{\theta}({\mathbf{x}})$ are generated by the same network $G_{\theta}$. Note that we only replace data distribution as $p^{(t_{k})}_{\theta}({\mathbf{x}})$ in the adversarial divergence for smoothing the learning objective, as recent work (Luo et al., 2024b) has found that it is beneficial to learn the generator with a mix of real data and revised data.
@@ 6.1 Text-to-Image Generation
Experiment Setting. We initialize the GAN generator by pre-trained PixArt-$\alpha$ which is a diffusion transformer with 0.6B parameters. For the GAN discriminator, we construct the latent discriminator by pre-trained SD 1.5 using the proposed construction introduced in Sec. 5.1. We switch the pre-trained PixArt-$\alpha$ to v-prediction by the proposed technique introduced in Sec. 5.1, followed by training on the JourneyDB dataset (Pan et al., 2023) with resizing as 512 resolution. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We observe the training convergence fast, requiring only 30k iterations and around 10 A800 days to be trained. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We also conduct experiments on fine-tuning SD 1.5 via LoRA to show the effectiveness of the proposed YOSO.
Figure 4: Qualitative comparisons of YOSO against competing methods. NFE denotes the Number of Function Evaluations.
@@
Table 2: Comparison of machine metrics on text-to-image generation across state-of-the-art methods. HFL denotes human feedback learning which might hack the machine metrics.
@@ 7 Related Work
Few-Step Text-to-Image Generation. LCM (Luo et al., 2023a) adapts consistency distillation (Song et al., 2023) for stable diffusion (SD) and enables image generation in 4 steps with acceptable quality. InstaFlow (Liu et al., 2023) adapts rectified flows (Liu et al., 2022; Liu, 2022) for SD, facilitating the generation of text-to-image in a mere single step. Despite this, the fidelity of the generated images is still poor. Some works (Yin et al., 2023) adapts variation score distillation (Wang et al., 2023b) and diff-instruct (Luo et al., 2023c) for SD, which requires training an additional DMs for the generator distribution, akin to the discriminator in GANs. We found models trained in this way seem to have serious mode collapse issues (see Fig. 5). Recently, combining DMs and GANs for one-step text-to-image generation has been explored. UFOGen (Xu et al., 2023c) extends DDGANs (Xiao et al., 2022) and SSIDMs (Xu et al., 2023b) to SD by modifying the computation of reconstruction loss from corrupted samples to clean samples. However, it still performs adversarial matching using corrupted samples. ADD (Sauer et al., 2023c) introduces a one-step text-to-image generation based on SD. It follows earlier research (Sauer et al., 2023a) by employing a pre-trained image encoder, DINOv2 (Oquab et al., 2024), as the backbone of the discriminator to accelerate the training. However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands. Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training. This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training. SDXL-Lightning (Lin et al., 2024b) combines progressive distillation and adversarial training in a progressive framework. Recently, a concurrent work Hyper-SD (Ren et al., 2024) combines consistency distillation and adversarial training in a progressive framework, while our work can be trained end-to-end. However, their work performs adversarial point-level matching via consistency loss, and they still perform adversarial training via injecting noise. This results in less effective one-step learning, thus they rely on human feedback learning for achieving better performance. In contrast, we perform the adversarial training at the distribution level by replacing real data with self-generated data to smooth the adversarial divergence, which not only can stabilize the training without injecting noise but also form effective one-step generation learning. Moreover, compared with above mentioned diffusion-GAN hybrid models, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them. Additionally, we extend our method not only to SD but also to PixArt-$\alpha$ (Chen et al., 2024) which is based on diffusion transformer (Peebles & Xie, 2022). This demonstrates the wide application of our proposed YOSO.
@@ 8 Conclusion
Acknowledgments
Jing Tang’s work is partially supported by National Key R&D Program of China under Grant No. 2023YFF0725100 and No. 2024YFA1012701, by the National Natural Science Foundation of China (NSFC) under Grant No. 62402410 and No. U22B2060, by Guangdong Provincial Project (No. 2023QN10X025), by Guangdong Basic and Applied Basic Research Foundation under Grant No. 2023A1515110131, by Guangzhou Municipal Science and Technology Bureau under Grant No. 2023A03J0667 and No. 2024A04J4454, by Guangzhou Municipal Education Bureau (No. 2024312263), and by Guangzhou Municipality Big Data Intelligence Key Lab (No. 2023A03J0012), Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628) and Guangzhou Municipal Key Laboratory of Financial Technology Cutting-Edge Research (No. 2024A03J0630).
@@ G.3 Ablation Study
Figure 9: Additional qualitative comparison on PixArt-$\alpha$ backbone at 1024 resolution.
Appendix H Additional Experiments
Table 5: Comparison of different methods of training from scratch on ImageNet-64.
[…]
BigGAN-deep 1 4.06
StyleGAN-XL 1 1.52
ADM 250 2.07
EDM 511 1.36
CT 1 13.0
iCT 1 3.25
YOSO (Ours) 1 2.65
Table 6: Comparison of different diffusion-gan hybrid methods on FFHQ-1024.
Method NFE FID ($\downarrow$)
Teacher Diffusion (Our reproduced) 250 13.88
Diffusion-GAN 1 22.16
Diffusion2GAN 1 18.01
UFOGen 1 24.37
YOSO (Ours) 1 16.23
Comparison to other Diffusion-Gan Hybrid Methods on High-Resolution Generation. We conducted comparative experiments between methods by fine-tuning pre-trained diffusion models. We trained a lightweight diffusion model in the latent space of DC-AE (Chen et al., 2025) for efficiency. Specifically, following DC-AE, we adopted DiT-S as the model backbone, which is a lightweight model with 33M parameters. For a fair comparison, we initialized the generator for YOSO as well as the other methods using the pre-trained diffusion model. As shown in Tab. 6, YOSO achieves better performance on FFHQ-1024 compared to other diffusion-gan hybrid distillation methods (Xu et al., 2023c; Wang et al., 2023a; Kang et al., 2024).
Training from scratch on ImageNet-64. To further demonstrate the ability of YOSO in training from scratch, we conducted additional experiments on ImageNet-64 using the EDM architecture, demonstrating YOSO’s effectiveness when trained from scratch. As shown in Tab. 5, our method achieves a better FID score compared to other methods, indicating both efficiency and improved performance.
Zero-Shot COCO FID. We argue that zero-shot COCO FID is not a good metric for evaluating text-to-image models. Existing work has found that zero-shot COCO FID can even be negatively correlated with human preferences in certain cases (Kirstain et al., 2023; Podell et al., 2023). We report the metric for completeness. We retrain the YOSO by fine-tuning on COYO-700M Byeon et al. (2022) and our internal dataset. As shown in Tab. 7, YOSO is capable of achieving a low zero-shot COCO FID-30k of 8.90, outperforming prior text-to-image GANs.
Model Latency $\downarrow$ FID-30k $\downarrow$ CLIP Score-30k $\uparrow$
StyleGAN-T 0.10s 13.90
GigaGAN 0.13s 9.09
SD v1.5 (CFG=7.5) 2.59s 13.45 0.32
SD v1.5 (CFG=3) 2.59s 8.78 0.32
LCM (4-step) 0.26s 23.62 0.30
Diffusion2GAN 0.09s 9.29
UFOGen 0.09s 12.78
InstaFlow-0.9B 0.09s 13.27 0.28
DMD 0.09s 14.93 0.32
SwiftBrush 0.09s 16.67 0.29
YOSO (full-finetuning) 0.09s 8.90 0.31
Table 7: Comparison of zero-shot COCO FID
Appendix I Additional Qualitative Comparison
We present an Additional Qualitative comparison to PixArt-$\alpha$ and PixArt-$\delta$ on 1024 resolution in Fig. 9. As can be observed, our method produces significantly better image quality compared to PixArt-$\delta$ (LCM) and achieves comparable results to the multi-step teacher model.
We present an Additional qualitative comparison among different sampling steps using YOSO-LoRA in Fig. 10. As can be observed, the 4-step samples have better visual quality than 1-step samples
We present an Additional qualitative comparison among 2-step examples of LCM by varying the distance metrics in Fig. 11 to verify the effectiveness of the proposed latent perceptual loss. In particular, we mainly vary the feature layer for computing the latent perceptual loss. It can be seen that using the bottleneck layer delivers better visual quality. More importantly, regardless of which layers are used to compute LPL, it is significantly better than MSE. This highlights the effectiveness of our proposed latent perceptual loss.
Figure 10: Qualitative comparison among different sampling steps using YOSO-LoRA.
Figure 11: Additional qualitative comparison of 2-step examples on varying the distance metric in LCM.